# The sphere complex of a locally finite graph

### v1

@@ 4 Short exact sequence
Lemmas 4.3 and 4.4 show that $\text{ker}(\Psi)$ is equal to $\text{Twists}(M_{\Gamma})$. To finish the proof, we need to show that if $[g]$ acts trivially on $\mathcal{S}(M_{\Gamma})$, then it lies in $\text{Twists}(M_{\Gamma})$.
The proof for when $\Gamma$ has rank 00 or rank at least $2$ follows from the proof of Lemma 4.3. Indeed, we may essentially repeat the proof as it did not rely on fixing $\pi_{2}(M_{\Gamma},x_{0})$, but instead fixing $\mathcal{S}(M_{\Gamma})$ (and spheres which bound a single puncture, which are also preserved as $g$ must clearly preserve the ends of $\Gamma$).
If $\Gamma$ has rank $1$ and at least $2$ ends, we may proceed as in Lemma 4.4 (Claim 4.5 isn’t needed) and decompose (a map isotopic to) $g$ into a map supported on the compact submanifold $K$, and maps supported on $M_{\Gamma}\setminus K$. We may assume $K$ has at least $2$ boundary components. On a component $U$ of $M_{\Gamma}\setminus K$, it follows from the previous paragraph that $g$ restricted to $U$ is isotopic to a composition of sphere twists. On $K$, we may further assume $g$ fixes the core sphere $S$ of $K$. But $K\setminus S$ is homeomorphic to $M_{0,s}$ for $s\geq 4$. Then $g$ cannot permute the boundary components of $K\setminus S$ as $g$ fixes $\mathcal{S}(K\setminus S)$, so by Proposition 3.3 and the discussion after it, $g$ restricted to $K$ is a composition of sphere twists. ∎
The assumption that in the rank $1$ case $\Gamma$ has at least $2$ ends only appears in this proof. If $\Gamma$ has no ends or $1$ end, then there is no guarantee that $g$ fixes the two boundary components of $K\setminus S$ corresponding to $S$, as $\mathcal{S}(K\setminus S)$ is empty.
@@ 5.1 Generating $\text{Twists}(M_{\Gamma})$
We will assume in this section that $\Gamma$ is in standard form. This implies that the boundary components of every piece are all separating spheres.
Recall the collection of spheres given in Lemma 4.3. This collection consists of the core spheres of the pieces of $M_{\Gamma}$, along with a collection of boundary spheres of pieces. We will show that we can remove the sphere twists on boundary components of pieces and still have a topological generating set of $\text{Twists}(M_{\Gamma})$.
@@
The primary tool used is the fact that we can move sphere twists along pairs of pants. Namely, if $P$ is an embedded copy of $S^{3}$ with three open balls removed in $M_{\Gamma}$, and $S_{1},S_{2}$, and $S_{3}$ the boundary components of $P$, then $T_{S_{1}}=T_{S_{2}}T_{S_{3}}$ in $\text{Map}(M_{\Gamma})$. This can be seen in Figure 2.1 of [hatcher_stabilization_2010].
@@
Corollary 5.2.
Let $\mathcal{S}$ be a (possibly infinite) proper collection of separating spheres in $M_{\Gamma}$. Let $f$ denote the product of all the sphere twists on $\mathcal{S}$. Then $[f]=\text{id}$ in $\text{Map}(M_{\Gamma})$.
Proof.
Every sphere twist $T_{S}$ on a sphere $S\in\mathcal{S}$ is isotopic to a composition of sphere twists on boundary components of pieces of $M_{\Gamma}$. To see this, note that $S$ lies in some minimal connected union of pieces, which we denote by $L$. Take $n$ and $s$ so that $L\cong M_{n,s}$. Then $S$ partitions $L$ into two pieces homeomorphic to $M_{n^{\prime},s^{\prime}}$ and $M_{n^{\prime\prime},s^{\prime\prime}}$ with $n^{\prime},n^{\prime\prime}\leq n$ and $s^{\prime},s^{\prime\prime}\leq s$. By applying moves similar to those in Figure 7, one can isotope $T_{S}$ so that it can be written as the product of sphere twists on the boundary (it doesn’t matter which component of $L\setminus S$ one pushes the sphere twists into).
We may thus isotope $f$ so that it is the product of sphere twists on the boundary sphere of pieces. To obtain this, apply the process in the previous paragraph for all the spheres in $\mathcal{S}$. The resulting concatenation of isotopies is well defined as $\mathcal{S}$ is a proper collection, and thus every piece of $M_{\Gamma}$ is only involved in finitely many isotopies. It follows from the discussion after Lemma 5.1 that $f=\text{id}$ in $\text{Map}(M_{\Gamma})$. ∎
In particular, we have the following.
@@
where $(f_{2})^{*}$ denotes the induced map on $H^{1}$ of $f_{2}$. The construction of this map does not rely on the compactness of the manifold, it relies only on the fact that there is a trivialization of the tangent bundle (which all oriented $3$-manifolds have). Thus we may define this map for any graph $\Gamma$.
@@
In particular, combining Lemmas 4.3, 4.4, 5.1, and 5.4, we obtain Theorem 1.2.
@@
The proofs of Lemmas 4.3 and 4.4 show that if $f\in\text{Diff}^{+}(M_{\Gamma})$ is homotopic to the identity, then it is isotopic to a product of sphere twists, and Lemma 5.1 specifies which sphere twists are necessary. It thus suffices to show that if such a product of sphere twists is homotopic to the identity, then it is isotopic to the identity as well.
@@ 5.2 Splitting the short exact sequence
Fix an open set in $\text{ker}(r)$, which we can write as $U\cap\text{ker}(r)$ for some $U\subset G$ open. It may not be the case that $\pi(U\cap\text{ker(r)})=\pi(U)$ (if it were, we would be done with this open set). Instead, we define another open set $U^{\prime}$ given by
@@
As $U$ is open and $r$ is continuous, it follows that $U^{\prime}$ is open. As $xr(x)^{-1}=x$ when $x\in\text{ker}(r)$, it follows that $U^{\prime}\cap\text{ker}(r)=U\cap\text{ker}(r)$. Further, for any $x\in U^{\prime}$, $\pi(x)=\pi(xr(x)^{-1})$ since $r(x)\in\text{ker}(\pi)$. But $xr(x)^{-1}\in U\cap\text{ker}(r)=U^{\prime}\cap\text{ker}(r)$, as by the definition of $U^{\prime}$, $xr(x)^{-1}\in U$, and
@@
where the second equality is using the fact that $r$ fixes $N$. Thus, $xr(x)^{-1}\in\text{ker}(r)$. This implies that $\pi(U^{\prime})=\pi(U^{\prime}\cap\text{ker}(r))$. Thus $\pi(U\cap\text{ker}(r))=\pi(U^{\prime}\cap\text{ker}(r))=\pi(U^{\prime})$ is open as $\pi$ is an open map. ∎
@@
By above, $\mathfrak{T}(g)\in U$. Since $g$ preserves $K$ and and $\mathfrak{T}(f)\in U$, it is clear that $g^{*}(\mathfrak{T}(f))\in U$. Hence $fg\in\mathfrak{T}^{-1}(U)$ and thus $f\mathcal{V}_{K}\subset\mathfrak{T}^{-1}(U)$. Since $f\in\mathfrak{T}^{-1}(U)$ is arbitrary, it follows that $\mathfrak{T}^{-1}(U)$ is open, implying $\mathfrak{T}$ is continuous.
The statement and argument in Lemma 5.4 along with Lemma 2.19 implies that $\mathfrak{T}$ restricts to a bijection from $\text{Twists}(M_{\Gamma})$ to $H^{1}(M_{\Gamma};{\mathbb{Z}}/2)$. Further, as $\text{Twists}(M_{\Gamma})$ acts trivially on $H^{1}(M_{\Gamma};{\mathbb{Z}}/2)$, this restriction is a homomorphism. Both groups are compact and Hausdorff, and hence the inverse map is also continuous.
@@ 6 Contractibility of the sphere complex and its subcomplexes
In this section, we consider the topological properties of $\mathcal{S}(M_{\Gamma})$ and some of its complexes. We note the following fact.
@@ 7.1 Topology
If $K$ is any compact subgraph of $\Gamma$, then take a graph $L$ containing $K$ so that $r^{-1}(L)$ is a union of pieces. Let $\mathcal{S}$ denote the finite collection of core spheres and boundary components of $L$. Then $\cap_{S\in\mathcal{S}}U_{S}\subset\mathcal{V}_{L}\subset\mathcal{V}_{K}$. Indeed, any lift to $\text{Map}(M_{\Gamma})$ of an element $f\in\cap_{S\in\mathcal{S}}U_{S}$ stabilizes $L$ up to isotopy, as the boundary spheres of $L$ are fixed, There is a well defined restriction of $f$ to $L$. But up to isotopy, as the core spheres of $L$ are also fixed, the restriction of any element in $\Psi^{-1}(f)$ to $L$ must be a product of sphere twists by Proposition 3.3, and thus $f\in\mathcal{V}_{L}$. ∎
We will now show that the topology on $\text{Map}(M_{\Gamma})$ generated by the $\mathcal{V}_{K}$ sets agrees with the quotient topology arising from $\text{Homeo}^{+}(M_{\Gamma})$ (which recall we can use instead of $\text{Diff}^{+}(M_{\Gamma})$), and also that this topology is Polish. We note that our argument here is inspired by the results in the appendix of [vlamis_notes_nodate].
@@
Fix a connected union of pieces $K\subset M_{\Gamma}$. If $\rho(f,\text{id})$ is sufficiently small, it follows that $\partial K$ and $f(\partial K)$ are homotopic. Thus by Theorem 4.1, we may modify $f$ by a small isotopy (supported in a small neighborhood of $\partial K$) so that $f$ fixes $\partial K$.
The mapping class group of $K$ is discrete. Indeed, this follows from work of Chernavskii or of Edwards–Kirby, which both say that homeomorphism groups of compact manifolds are locally contractible, and thus locally path connected [chernavskii_local_1969][edwards_deformations_1971]. Hence the image in $\mathrm{Map}(K)$ of a small metric ball in $\text{Homeo}(K,\partial K)$ is a point. Letting $\pi$ denote the quotient map from $\text{Homeo}^{+}(M_{\Gamma})$ to $\text{Map}(M_{\Gamma})$, it follows that for any $[f]\in\mathcal{V}_{K}$, the $\pi$ image of a small $\rho$-ball about $f$ is an open set containing $[f]$ which is contained in $\mathcal{V}_{K}$.
On the other hand, the $\pi$ image of a small open ball about $f\in\text{Homeo}^{+}(M_{\Gamma})$ clearly contains $[f]\mathcal{V}_{K}$ for $K$ a sufficiently large union of pieces. This is by the definition of $\rho$. Thus the topologies are equivalent. ∎
@@
This follows from Theorem 1.1, as $\text{Map}(M_{\Gamma})$ is topologically the product of two Polish spaces (in either of the two topologies, which are equivalent by Proposition 7.2), and is thus Polish. As $\text{PMap}(M_{\Gamma})$ is a closed subgroup of $\text{Map}(M_{\Gamma})$, it is Polish because closed subgroups of Polish groups are Polish. ∎

### v2

@@ 4 Short exact sequence
Lemmas 4.3 and 4.4 show that $\text{ker}(\Psi)$ is equal to $\text{Twists}(M_{\Gamma})$ (regardless of if $\Gamma$ is sporadic or not). Now assume that $\Gamma$ is non-sporadic. To finish, we need to show that if $[g]\in\text{Map}(M_{\Gamma})$ acts trivially on $\mathcal{S}(M_{\Gamma})$, then it lies in $\text{Twists}(M_{\Gamma})$.
Fix a compact exhaustion of $M_{\Gamma}$ by connected unions of pieces. By assumption, $g$ fixes the isotopy classes of the core spheres of $M_{\Gamma}$ as well as the nonperipheral boundary components of the pieces. Lemma 4.6 shows that the isotopy classes of peripheral boundary spheres of pieces are also fixed. Thus, the proof of Lemma 4.3 shows that, for any graph $\Gamma$, that there is a $f\in[g]$ which stabilizes the core spheres and the nonperipheral boundary spheres of pieces of $M_{\Gamma}$.
We are not assuming $g$ fixes isotopy classes of embeddings of spheres. As $g$ is orientation preserving, either $f$ fixes a given sphere pointwise (up to isotopy), or the diffeomorphism $\bar{f}$ induced from $f$ by cutting along each of these spheres permutes a pair of spheres obtained from cutting along a single sphere in the collection above. This happens if $f$ flips the co-orientation of the sphere in question. If we show these sorts of permutations don’t happen, then we will be done, as the restriction of $\bar{f}$ to each component will be a product of sphere twists, again by Lemma 2.5 of [brendle_mapping_2023] or Proposition 3.3.
Suppose first that $S$ is a separating sphere so that $\bar{f}$ permutes the two boundary components corresponding to $S$. Let $U$ and $U^{\prime}$ denote the two components of $M_{\Gamma}\setminus S$. Then $f$ swaps $U$ and $U^{\prime}$ by the assumptions on $\bar{f}$ and $S$. If $U$ contains an element $S^{\prime}\in\mathcal{S}(M_{\Gamma})$ not equal to $S$, then $S^{\prime}\neq g(S^{\prime})$. But this is impossible by the assumption on $g$, so no such $S^{\prime}$ exists. This can only happen if $U$ and $U^{\prime}$ have empty sphere graphs, which implies $M_{\Gamma}$ is a sphere with at most $4$ punctures. Hence $\Gamma$ is sporadic, a contradiction.
Now suppose $S$ is nonseparating so that $\bar{f}$ permutes the two boundary components corresponding to $S$. Hence $\Gamma$ must have rank at least $1$. As $\Gamma$ is non-sporadic by assumption, we can take a diffeomorphic copy $K$ of $M_{1,2}$ embedded in $M_{\Gamma}$ containing $S$ as a nonseparating sphere and so that the boundary spheres of $K$ are essential and separating (but potentially peripheral) in $M_{\Gamma}$. By the above discussion about separating spheres and the assumption on $\partial K$, $f$ restricts (up to isotopy) to a diffeomorphism on $K$ fixing $\partial K$ and stabilizing $S$. But if the diffeomorphism induced on $K\setminus S$ swapped the two boundary components coming from $S$, then it also swaps the two nonperipheral spheres in $K\setminus S\cong M_{0,4}$ that separate the two boundary components coming from $S$. Thus $f$ itself could not fix these two spheres, so we again have a contradiction. The result follows.
We now detail why the sporadic cases are excluded. If $\Gamma$ is rank 00 with at most $3$ ends, then $\mathcal{S}(M_{\Gamma})$ is empty and the result is vacuous. If $\Gamma$ is a rank 00 graph with $4$ ends, then elements of $\text{Map}(\Gamma)$ which correspond to products of disjoint cycles acting on the ends of $\Gamma$ fix all three elements of $\mathcal{S}(M_{\Gamma})$. In other words, the map $\text{Map}(\Gamma)\to\text{Aut}(\mathcal{S}(M_{\Gamma}))$ is just the exceptional homomorphism $S_{4}\to S_{3}$. If $\Gamma$ is rank $1$ with at most $1$ end, then $\mathcal{S}(M_{\Gamma})$ is a single vertex, and $\text{Map}(\Gamma)\cong{\mathbb{Z}}/2$ and thus there is trivially not a faithful action. Finally, if $\Gamma$ is a compact rank $2$ graph, then $\text{Map}(\Gamma)\cong\text{Out}(F_{2})\cong GL(2,{\mathbb{Z}})$ is well known to surject onto $\text{Aut}(\mathcal{S}(M_{\Gamma}))\cong\text{PGL}(2,{\mathbb{Z}})$ with an order $2$ kernel generated by the automorphism that flips the signs of a fixed generating of $F_{2}$. In particular, the action is not faithful.
@@ 5.1 Generating $\text{Twists}(M_{\Gamma})$
We again assume in this section that $\Gamma$ is in standard form and that $\Gamma$ has no vertices of valence $1$. As discussed before, this implies that the boundary components of every piece are all separating spheres.
Recall the collection of spheres given in Lemmas 4.3 and 4.4. This collection consists of the core spheres of the pieces of $M_{\Gamma}$, along with a collection of boundary spheres of pieces. By the two lemmas, every element of $\text{Twists}(M_{\Gamma})$ can be written as a product of sphere twists on these spheres. The next lemma shows that we can remove the sphere twists on boundary components of pieces.
@@
The primary fact needed is that we can move sphere twists along pairs of pants. Namely, if $P$ is an embedded copy of $S^{3}$ with three open balls removed in $M_{\Gamma}$, and $S_{1},S_{2}$, and $S_{3}$ the boundary components of $P$, then $T_{S_{1}}=T_{S_{2}}T_{S_{3}}$ in $\text{Map}(M_{\Gamma})$. This can be seen in Figure 2.1 of [hatcher_stabilization_2010].
@@
where $(f_{2})^{*}$ denotes the induced map on $H^{1}$ of $f_{2}$. The construction of this map does not rely on the compactness of the manifold, it relies only on the fact that there is a trivialization of the tangent bundle (which all oriented $3$-manifolds have, see Theorem 1 in Section VII of [kirby_topology_1989]). Thus we may define this map for any graph $\Gamma$.
@@
Proof of Theorem 1.2.
Since $\text{Twists}(M_{\Gamma})\subset\text{ker}(\Psi)$, Lemmas 4.3 and 4.4 in particular imply that any element of $\text{Twists}(M_{\Gamma})$ can be written as a product of sphere twists on the core spheres, $\mathcal{C}$, and boundary components of pieces of $M_{\Gamma}$. Lemma 5.1 shows that the sphere twists on boundary components of pieces are trivial, and Lemma 5.3 implies that any nontrivial composition of sphere twists on the core spheres of $M_{\Gamma}$ is not isotopic to the identity.
Fixing a bijection $\alpha:\mathcal{C}\to\{1,\ldots,\text{rk}(\Gamma)\}$ (allowing for $\text{rk}(\Gamma)=\infty$, in which it is not included in the set), we obtain an induced isomorphism $\tau:\text{Twists}(M_{\Gamma})\to\Pi_{i=1}^{\text{rk}(\Gamma)}{\mathbb{Z}}/2$. The map $\tau$ is obtained by sending a composition of sphere twists on $\mathcal{C}^{\prime}\subset\mathcal{C}$ to the tuple $(a_{i})_{i=1}^{\text{rk}(\Gamma)}$ with $a_{i}=1$ if $\alpha^{-1}(i)\in\mathcal{C}^{\prime}$, and $a_{i}=0$ otherwise.
We wish to show that $\tau$ is also a homeomorphism. Fix a finite union of pieces $K$. Let $U_{K}\leq\Pi_{i=1}^{\text{rk}(\Gamma)}{\mathbb{Z}}/2$ be the subgroup of tuples so that the $i$’th component is 00 if $\alpha^{-1}(i)$ lies in $K$. The collection of all $U_{K}$’s is a basis at the identity, so to show that $\tau$ is a homeomorphism it suffices to check that $\tau(\mathcal{V}_{K}\cap\text{Twists}(M_{\Gamma}))=U_{K}$, i.e. $\tau$ sends a basis to a basis. To see this, note that an element $f\in\text{Twists}(M_{\Gamma})$ lies in $\mathcal{V}_{K}$ precisely when no sphere twist $T_{S}$ that appears in $f$ is so that $S\subset K$. This is because, for a core sphere $S^{\prime}\subset K$, $T_{S^{\prime}}$ cannot be isotoped so that it fixes $K$, as $\mathfrak{T}([T_{S^{\prime}}])$ would then evaluate to 00 on any $\gamma\in H_{1}(M_{\Gamma};{\mathbb{Z}}/2)$ which lies in $K$. But this is clearly contradicts the proof of Lemma 5.3 as one can can choose the $\gamma$ in that proof to lie in $K$. Hence $\tau$ sends elements of $\mathcal{V}_{K}\cap\text{Twists}(M_{\Gamma})$ to $U_{K}$. Further, it is easy to see that $\tau(\mathcal{V}_{K}\cap\text{Twists}(M_{\Gamma}))$ contains $U_{K}$ by the definition of $\tau$, finishing the proof. ∎
@@
The proof of Theorem 1.1(1) shows that if $f\in\text{Diff}^{+}(M_{\Gamma})$ is homotopic to the identity, then it is isotopic to a product of sphere twists, and Lemma 5.1 specifies which sphere twists are necessary. It thus suffices to show that if such a product of sphere twists is homotopic to the identity, then it is isotopic to the identity as well.
@@ 5.2 Splitting the short exact sequence
Fix an open set in $\text{ker}(r)$, which we can write as $U\cap\text{ker}(r)$ for some $U\subset G$ open. It may not be the case that $\pi(U\cap\text{ker(r)})=\pi(U)$ (if it were, we would be done with this open set as $\pi$ is an open map). Instead, we define another open set $U^{\prime}$ given by
@@
As $U$ is open and $r$ is continuous, it follows that $U^{\prime}$ is open. As $xr(x)^{-1}=x$ when $x\in\text{ker}(r)$, it follows that $U^{\prime}\cap\text{ker}(r)=U\cap\text{ker}(r)$.
Fix $x\in U^{\prime}$. Then $xr(x)^{-1}\in U\cap\text{ker}(r)=U^{\prime}\cap\text{ker}(r)$, as by the definition of $U^{\prime}$, $xr(x)^{-1}\in U$, and
@@
where the second equality is using the fact that $r$ fixes $N$. Thus, $xr(x)^{-1}\in\text{ker}(r)$. Also, $\pi(x)=\pi(xr(x)^{-1})$ since $r(x)\in\text{ker}(\pi)$. This implies that $\pi(U^{\prime})=\pi(U^{\prime}\cap\text{ker}(r))$. Thus $\pi(U\cap\text{ker}(r))=\pi(U^{\prime}\cap\text{ker}(r))=\pi(U^{\prime})$ is open as $\pi$ is an open map. ∎
@@
By above, $\mathfrak{T}(g)\in U$. Since $g$ preserves $K$ and and $\mathfrak{T}(f)\in U$, it is clear that $g^{*}(\mathfrak{T}(f))\in U$. This is because $g$ stabilizes the homology classes of both the spheres in $K$ and in $M_{\Gamma}\setminus K$. Hence $fg\in\mathfrak{T}^{-1}(U)$ and thus $f\mathcal{V}_{K}\subset\mathfrak{T}^{-1}(U)$. Since $f\in\mathfrak{T}^{-1}(U)$ is arbitrary, it follows that $\mathfrak{T}^{-1}(U)$ is open, as it is the union of open sets. Thus $\mathfrak{T}$ is continuous.
The statement and argument in Lemma 5.3 along with Lemma 2.19 implies that $\mathfrak{T}$ restricts to a bijection from $\text{Twists}(M_{\Gamma})$ to $H^{1}(M_{\Gamma};{\mathbb{Z}}/2)$. Further, as $\text{Twists}(M_{\Gamma})$ acts trivially on $H^{1}(M_{\Gamma};{\mathbb{Z}}/2)$, this restriction is a homomorphism. Both groups are compact and Hausdorff, and hence the inverse map is also continuous. (We remark also that the restriction of $\mathfrak{T}$ to $\text{Twists}(M_{\Gamma})$ agrees with the map $\tau$ defined in the proof of Theorem 1.2, and there it is shown that $\tau$ is a homeomorphism as well).
@@ 6 Contractibility of the sphere complex and its subcomplexes
In this section, we consider the topological properties of $\mathcal{S}(M_{\Gamma})$ and some of its subcomplexes. We note the following fact.
@@ 7.1 Topology
If $K$ is any compact subgraph of $\Gamma$, then take a graph $L$ containing $K$ so that $r^{-1}(L)$ is a union of pieces. Let $\mathcal{S}$ denote the finite collection of core spheres and boundary components of $L$. Then $\cap_{S\in\mathcal{S}}U_{S}\subset\mathcal{V}_{L}\subset\mathcal{V}_{K}$. Indeed, any lift to $\text{Map}(M_{\Gamma})$ of an element $f\in\cap_{S\in\mathcal{S}}U_{S}$ stabilizes $L$ up to isotopy, as the core spheres and boundary spheres of $L$ are fixed, There is a well defined restriction of $f$ to $L$. But up to isotopy, as the core spheres of $L$ are also fixed, the restriction of any element in $\Psi^{-1}(f)$ to $L$ must be a product of sphere twists by Proposition 3.3, and thus $f\in\mathcal{V}_{L}$. ∎
We will now show that the topology on $\text{Map}(M_{\Gamma})$ generated by the $\mathcal{V}_{K}$ sets agrees with the quotient topology arising from $\text{Homeo}^{+}(M_{\Gamma})$ (which recall we can use instead of $\text{Diff}^{+}(M_{\Gamma})$), and also that this topology is Polish. The argument here is inspired by the results in the appendix of [vlamis_notes_nodate].
@@
Let $\pi$ denote the quotient map from $\text{Homeo}^{+}(M_{\Gamma})$ to $\text{Map}(M_{\Gamma})$. Fix a connected union of pieces $K\subset M_{\Gamma}$. If $\rho(f,\text{id})$ is sufficiently small then $\partial K$ and $f(\partial K)$, are homotopic (as embeddings). Thus by Theorem 4.1, we may modify $f$ by a small isotopy so that $f$ fixes $\partial K$ and also stabilizes $K$. But the group $\text{Map}(K,\partial K)$ is discrete. Indeed, this follows from work of Chernavskii or of Edwards–Kirby, which both say that homeomorphism groups of compact manifolds are locally contractible, and thus locally path connected [chernavskii_local_1969][edwards_deformations_1971]. As $f$ restricted to $K$ is close to the identity on $K$ (assuming $\rho(f,id)$ is sufficiently small), it follows that $\pi(f)\in\mathcal{V}_{K}$.
Note that $\pi$ is an open map as quotient homomorphisms of topological groups are open. It follows that for any $[f]\in\mathcal{V}_{K}$, the $\pi$ image of a small $\rho$-ball about a representative $f$ of $[f]$ fixing $K$ is an open set (in the quotient topology) containing $[f]$ which is also a subset of $\mathcal{V}_{K}$. Indeed, if $\rho(f,f^{\prime})$ is small, then since $f$ is assumed to fix $K$ pointwise, the work in the previous paragraph shows that $f^{\prime}$ can be isotoped to fix $K$ pointwise as well. Thus $\mathcal{V}_{K}$ is a union of open sets in the quotient topology, so it is itself open.
On the other hand, for any open set $U$ in the quotient topology and any $[g]\in U$ there is some sufficiently large union of pieces $K$ so that $[g]\mathcal{V}_{K}\subset U$. This is because homeomorphisms in $\pi^{-1}([g]\mathcal{V}_{K})$ have isotopy representatives whose restrictions to $K$ are the same as that of $g$, a fixed representative of $[g]$. In particular, by the definition of the metric $\rho$, every element of $\pi^{-1}([g]\mathcal{V}_{K})$ has a representative $g^{\prime}$ so that $\rho(g,g^{\prime})<\epsilon$, where $\epsilon>0$ is small enough so that the $\pi$ image of the $\epsilon$ $\rho$-ball about $g$ lies in $U$ (allowing $K$ to be as large as desired is what allows us to guarantee this for any $\epsilon$). Thus $U$ is the union of open sets in the topology generated by the collection of $\mathcal{V}_{K}$’s, and is thus open in that topology as well. It follows that the topologies are equivalent. ∎
@@
This follows from Theorem 1.1, as $\text{Map}(M_{\Gamma})$ is topologically the product of two Polish spaces, and is thus Polish. As $\text{PMap}(M_{\Gamma})$ is a closed subgroup of $\text{Map}(M_{\Gamma})$, it is Polish because closed subgroups of Polish groups are Polish. ∎